A man moving down in inclined plane feels the rain to come vertically downward. The speed of man is 2root3 and the angle of inclination is 30 then find the angle at which he should kept his umbrella to avoid rain on itself and the actual speed of rain is 5m/s

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
* cos30 = 5 cosa ( a =angle of rain from horizontal)cosa = 3/5, a = 53 degressWhen man is at rest, he should hold his umbrella in the actual direction of rain, which is 53 degress from horizontal.\sqrt{3}If you draw the vector diagram of the velocities when the man is in motion, the resultant direcion of the rain is vertical.Horizontal component of the velocity of man should cancel with that of the rain to make the resultant vertical.2 RegardsArun (askIITians forum expert)